A taxi cab company has 12 Ambassadors and 8 Fiats. If 5 of these taxi cabs are in the shop for repairs and Ambassador is as likely to be in for repairs as a fiat. what is the probability that (i) 3 of them are Ambassadors and 2 are fiats? (ii) at least 3 of them are Ambassadors? And (iii) all 5 of them are of the same make?


1
Expert's answer
2021-07-19T05:50:15-0400

"12+8=20"

(i)


"P(3A \ \&\ 2F)=\dfrac{\dbinom{12}{3}\dbinom{8}{2}}{\dbinom{20}{5}}=\dfrac{220(28)}{15504}"

"=\dfrac{385}{969}\approx0.3973"

(ii)


"P(at\ least\ 3A)=\dfrac{\dbinom{12}{3}\dbinom{8}{2}}{\dbinom{20}{5}}+\dfrac{\dbinom{12}{4}\dbinom{8}{1}}{\dbinom{20}{5}}"

"+\dfrac{\dbinom{12}{5}\dbinom{8}{0}}{\dbinom{20}{5}}=\dfrac{220(28)+495(8)+792(1)}{15504}""=\dfrac{682}{969}\approx0.7038"

(iii)


"P(5\ same)=\dfrac{\dbinom{12}{5}\dbinom{8}{0}}{\dbinom{20}{5}}+\dfrac{\dbinom{12}{0}\dbinom{8}{5}}{\dbinom{20}{5}}"

"=\dfrac{792(1)+1(56)}{15504}=\dfrac{53}{969}\approx0.0547"